The following data are available for two divisions of Solomons Company:

North Division South Division
Division operating profit $ 8,695,000 $ 45,780,000
Division investment 37,000,000 327,000,000

The cost of capital for the company is 8 percent. Ignore taxes.  

Required:

a-1. Calculate the ROI for both North and South divisions. (Enter your answers as a percentage rounded to 1 decimal place (i.e., 32.1).)

a-2. If Solomons measures performance using ROI, which division had the better performance?

North
South

b-1. Calculate the EVA for both North and South divisions. (The divisions have no current liabilities.)

b-2. If Solomons measures performance using economic value added, which division had the better performance?

North
South

c. Would your evaluation change if the company’s cost of capital were 19.0 percent?

1. When evaluated by ROI?

Yes
No

2. When evaluated by EVA?

Yes
No
Answer #1

a-1 ROI = Operating profit/Investment

North = $8695000/37000000 * 100 = 23.5%

South = 45780000/327000000 * 100 = 14.0%

a-2 The division with higher ROI is better.

Hence, North division is performing better.

b-1 EVA = Operating profit - (Investment* required rate)

North = $8695000 - ($37000000*8%)

= $5735000

South = $45780000 - ($327000000*8%)

= $19620000

b-2 The division with higher EVA is performing better.

Hence according to EVA South division is performing better

c. 1. No, ROI shall remain the same

2. at 19%

EVA of north division = 8695000 - (37000000*19%)

= $1665000

EVA of south division = 45780000 - (327000000*16%)

= -$6540000

Yes, now north division is better performing.